The wedding dress that liels wanted to purchase,cost R15000.she purchase it using her credit card or she could pay it off over 3 uears if she made use of hire purchase agreement.


1.determine how much liels would eventually pay if she decieded to use her credit card where the rate was 18% compound interest per annum over 3 years.


2.determine how much liels would eventually pay if she decieded to the hire purchase agreement at a rate of 18% per annum simple interest over 3 years


3.The original cash price for the dress was R15000.in 3 uears time,due to inflaction,the cost of the same dress would be R18017.36

What would be the average inflation rate(as a percentage) for the period?


1
Expert's answer
2021-03-24T13:49:10-0400

1.

"A=P(1+r\/n)^nt"

"15000=P(1+0.18)^3"

"15000=1.643032P"

"P=9129.46309"

Liel would pay R9129.46309 if she paid using her credit card.


2.

"cash price = 9129.46309"

"SI= P*R*T"

"= 9129.46309*0.18*3"

"=4929.910069"

"total amoumt =9129.46309+4929.910069"

"=R14059.37316"


3.

"[(18017.36-15000)\/15000]*100"

"=0.201157*100"

"=20.1157%" percent